# A Survey of Conversational Search

### v1

@@ 3.1.1 Available Datasets
There are several available datasets for clarification in conversational retrieval with different impacts, which can be used to support the development of methodologies. Qulac [19] and ClariQ [179] are two commonly used datasets that are built on top of the TREC Web Track 2009-2012 data [204]. Since these two datasets only contain clarifying questions and answers written by humans by crowd sourcing which are not applicable for many scenarios e.g., training a generation model, Aliannejadi et al. [205] further collect synthetic conversations to enhance these two and benchmarks several state-of-the-art methods for search clarification. They also propose a pipeline for online and offline evaluation.  Sekulic et al. [84] propose another incremental dataset ClariQ-FKw based on ClariQ. They extracted facet keywords in clarifying questions as the key information. Recently, Yuan et al. [87] release the first multi-modal clarifying question dataset Melon which extends existing study scenarios by incorporating multi-modal information. The construction of the datasets for search clarification in conversational retrieval is crucial and makes the development of the systems possible.
@@ 3.2 Web Search Clarification
The aforementioned studies for search clarification in conversational retrieval are limited by only selecting the clarifying questions and evaluating the response in a fixed set. Different from them, Zamani et al. [88] propose clarification in web search, In web search, the form of the user’s query is more flexible, which can be a conversational query or a keyword query. When the user’s query is ambiguous or faceted, the system can ask the user a clarification question, and provide multiple candidate facets for the user to choose from, which form a clarification pane as shown in Figure 5 (b). In the clarification pane, the clarification question is usually generated to improve the user’s search experience, and the candidate facets are a set of generated or selected terms representing a group of high-coherency user sub-intents that the user may be interested in. When the user clicks one of the given facets, the query will be reconstructed by concatenating with the selected facet to form a new query and retrieve a new list of web pages [91].
@@ 3.2.3 Generating Candidate Facets
Unlike search clarification in other scenarios, web search clarification not only asks a clarifying question but also provides users with several candidate facets to click on. This facilitates user interaction and is more suitable for open-domain scenarios. Before this, query facets is a common query intent mining solution, aiming to mine comprehensive, multi-dimensional query sub-intents. For example, there are some rule-based and machine learning-based methods to mining these terms from the search result pages of the query [93, 94, 95].
Different from these studies, web search clarification focuses on finding a set of facets that users are most interested in, rather than showing all of them to users, which helps to concentrate on users’ interests and provides a better user experience. Following this idea, Hashemi et al. [96, 97] first propose generating these facets using a seq2seq framework.  Samarinas et al. [98] summarize all rule-based, machine learning-based, and pre-trained language model-based mainstream methods and models to generate the candidate facets for search clarification and conclude that BART-based models taking the query and search result snippets as the input show the best performance in almost all evaluation metrics. Zhao et al. [92] deem that candidate facets are usually in the form of lists, so they collect parallel structures from the search result pages and use this information to enhance the generation ability of the BART-based model. Litvinov et al. [99] analyze the whether and how the coherency of candidate facets influence the generation performance and proposed a coherency-based method for generation. Zhao and Dou [100] think that direct generation of candidate facets is not suitable in a multi-turn manner, so they propose another framework to generate candidate facets in a multi-turn manner.
Recently, with the development of LLMs, some LLM-based methods have been proposed for generating candidate facets for web search clarification. Ni et al. [101] conduct a comparative study to compare different training objectives and base models for generating clarification facets. Lee and Kim [102] fully rely on LLMs to generate the facets. We believe that facets generation based on LLMs will become a more effective approach because LLMs contain much general and domain-specific knowledge.
@@ 3.3 Search Clarification for Question Answering
In question answering (QA), the goal is to answer the user’s specific question. However, the question can contain ambiguity, especially under a conversational scenario. Thus, it is necessary to ask a question for disambiguation. In this section, we classify existing studies in search clarification for QA as (1) Community Question Answering and (2) Knowledge-based Question Answering. As shown in Figure 5 (c), for Community QA, the system usually focuses on ambiguous intents in a domain-specific scenario such as programming, while the clarification usually focuses on some entity ambiguities in the knowledge-based QA.
@@ 3.3.1 Datasets
In some community QA studies [107, 108], the data is extracted from some community QA websites such as StackExchange and Amazon. These websites contain lots of ambiguous questions, user answers, and clarifying questions. Later, Kumar and Black [178] propose a dataset ClarQ based on StackExchange. To our best knowledge, the scale of ClarQ is much larger than other datasets based on StackExchange. On the other hand, in knowledge-based question answering, Abg-CoQA [182] is proposed to tackle ambiguity in conversational QA which focuses on disambiguation for human-to-human conversations.
@@ 3.3.2 Community Question Answering
Braslavski et al. [208] make the first step to analyze clarifying questions in community QA. They study the user questions from two domains and analyze the user behavior and the taxonomy of the asked clarification questions. After that, Rao et al. [107, 108] design different methods to generate good clarifying questions for community QA. Trienes and Balog [109] is the first study investigating whether a community QA question is clear or not. They propose a method of judging whether a question needs further clarification. Kumar et al. [110] study how to rank clarifying question candidates by natural language inference. As the clarification mechanism is complicated in community QA, some studies [209, 210] made some analytical studies about the feature of good clarifying questions, the taxonomy of clarification questions, and the types of user answers. Although these studies have proposed effective methods to generate community QA questions, these methods are difficult to transfer to other open-domain scenarios.
@@ 3.3.3 Knowledge-based Question Answering
In knowledge-based question answering systems, the clarification process usually targets reference issues between entities. Xu et al. [111] propose a dataset CLAQUA focus on solving clarification in knowledge-based question answering. They also figure out single-turn and multi-turn ambiguity and design to tackle these two situations. Nakano et al. [112] explore the impact of generating pseudo-ambiguous and clarifying questions based on sentence structure on the clarification question answering system. This study provides useful guidance for optimizing the clarification question answering system, enabling it to better understand user needs and generate more targeted clarification questions. Kim et al. [113] study the method of using retrieval-augmented large language models to handle ambiguous questions. Lee et al. [20] provide an in-depth analysis for understanding how to interact with users in open-domain question answering systems and how to adjust their behavior according to user needs. The above studies have solved the entity ambiguity problem effectively, which is important for understanding user’s ambiguous intent in conversational search.
@@ 3.4 Domain-specific Search Clarification
Aside for the aforementioned general search clarification scenarios, some studies focus on a closed search scenario like Conversational Recommender System, Legal Search, etc, are also important. Figure 5 (d) illustrates two close-domain settings, which require the methods should be accordingly adapted to the scenarios. We introduce and summarize related studies for different domain-specific search clarification scenarios in this section.
@@ 3.4.2 Legal Search Clarification
In legal search, users often deliver ambiguous intents due to the complexity of legal documents. Therefore, a clarifying question can be effective in helping the user articulate her complicated search intent. Liu et al. [49] compare the search behavior and outcomes of conversational and traditional search in legal case retrieval. After that, they further explore how buffer mechanisms and generated queries can help improve the performance of conversational agents in legal case retrieval [144]. Furthermore, they also try to study the behavior of conversational agents for retrieving better legal cases. This study provided an in-depth analysis for understanding how conversational agents interact with users and how to adjust their behavior according to user needs. Recently, Liu et al. [226] introduce event schema. By applying event schema, conversational agents can have a better understanding of the user’s potential intent and ask targeted clarifying questions. Therefore, the accuracy and efficiency of retrieving legal cases can be improved.
@@ 3.4.3 Search Clarification for Other Scenarios
Search clarification can further be applied in any scenario where the user’s query or question can be ambiguous or faceted. In this section, we list some other scenarios of search clarification. These studies are used to expand clarification to new scenarios and can provide inspiration.
  • Interactive classification [88]
  • Image guessing [227]
  • Product description [228, 229]
  • Collaborative building [230]
  • Programming [231, 232]
  • User simulation [233]
  • Social or moral situation [11]
  • Task-oriented dialogues [21]
  • Math [234]
  • Spatial reasoning [235]
  • Collaborative dialogue [236]
@@ 3.6 Future Discussion
Search Clarification for LLMs In Section 3.5, we summarize recent studies about search clarification with LLMs. These studies focus on utilizing the ability of LLMs as a tool for improving search clarification effectiveness in conversational search systems. However, in some cases, the LLM itself can also be seen as a conversational search system that has a large amount of user logs. Search clarification for LLMs themselves is also worth studying.
Search Clarification for Retrieval-Augmented Generation Some current hot search scenarios can also introduce clarification for enhancement, such as Retrieval-Augmented Generation (RAG). For example, when the user asks a question including an ambiguous entity, the system can ask a question to let the user disambiguate the query and then do the retrieval process. The system can also identify ambiguous or faceted intents embodied in the question to ask more natural clarifying questions.
Search Clarification in Vertical Domains We can explore more natural clarification generation in some vertical fields, such as medical consultation, customer service robots, etc.
@@ 4 Conversational Retrieval
First, although recent neural models, e.g., BERT [237], T5 [238], and recently emerged LLMs [239], have shown the advanced ability of sequence modeling, it is still a long-term challenge to fully interpret the multi-turn natural language form conversations. Second, the transitions of search intents will introduce much noise into the historical context, which is irrelevant to the current search behavior [24, 62]. Thus, it is crucial to address how to denoise the conversation context. Third, the training data for conversational retrieval models are limited. Labeling multi-turn interactions is time-consuming for human annotators and the relevance signals of passages are difficult to gain. Therefore, various levels of data augmentation techniques, such as generating queries, context-passage pairs, and whole conversations have been developed. Lastly, though conversational retrieval models can usually achieve better performance than query reformulation approaches, the lack of interpretability hampers the understanding of dense representations in conversations. To address these challenges, the existing studies of conversational retrieval can be categorized into four groups: (1) conversation modeling, (2) context denoising, (3) data augmentation, and (4) interpretation.
@@ 4.1 Conversation Modeling
Figure 6: An example illustrating the end-to-end learning of conversational retrieval.
@@ 4.2 Context Denoising
Within a conversational session, not all historical query turns are relevant to address the current user information needs. Simply concatenating all historical context would inject unnecessary and noisy turns and thus degrade the performance of the system. How to accurately identify the useful historical context of the conversation is a crucial research question.
@@ 4.2.1 Implicit Context Denoising
Some works attempt to perform context denoising implicitly. For example, Yu et al. [122] leverages dense retrieval to embed conversational queries and documents into a shared vector space where semantic similarity can be directly measured. It employs a teacher-student framework, which involves using an existing, well-trained dense retriever as a teacher to guide the student model, by inheriting its document encodings and learning to mimic the teacher’s embeddings for queries. This strategy enhances model ability by capturing the necessary context while ignoring unrelated information to improve retrieval accuracy across conversational turns. Mao et al. [24] develop a framework by combining curriculum learning and contrastive learning to progressively guide a conversational query encoder on how to filter out irrelevant information from the conversation history. It involves generating noisy and clean versions of conversational data and using these to train the model by gradually increasing the complexity of the training data. Krasakis et al. [123] aim to address context denoising by leveraging existing ad-hoc search data to understand and filter out noisy context, improving the relevance judgment of retrieved passages without additional fine-tuning on conversational datasets. Specifically, it contextualizes the query embeddings within the conversation while ensuring that only the current query terms are matched with potential answers. More recently, Mo et al. [124] propose QRACDR to first define a better aligned embedding space for the query representation learning and enable to model to learn to align the query representation with this area via rewritten queries and relevance judgments, which obtains a less noisy context representation implicitly.
@@ 4.3.2 Contrastive Sample Generation
Sufficient supervision signals, either positive or negative, are important for dense retrieval [241] with contrastive learning [242]. In terms of conversation scenarios, additional contrastive samples can be generated from the conversational history. The basic assumption [27] is among the top-ranked historical retrieved results, some of them can be highly relevant to the current query, which can be used as pseudo positive samples, while others are less relevant and can serve as hard negatives for training. Besides, Kim et al. [127] identify a “retrieval shortcut” phenomenon, which occurs when models overly rely on partial conversation history, but ignore the current query and results in topic drift. To this end, this work mines additional hard negatives from the historical turns to mitigate the dependency on retrieval shortcuts by forcing the model to consider both the current question and the entire conversation history.

### v2

@@ 3.1.1 Available Datasets
Several datasets for conversational retrieval clarification are available to support methodology development. Qulac [21] and ClariQ [209] are two commonly used datasets that are built on top of the TREC Web Track 2009-2012 data [238]. Since these two datasets only contain clarifying questions and answers written by humans by crowd sourcing which are not applicable for many scenarios e.g., training a generation model, Aliannejadi et al. [239] further collect synthetic conversations to enhance these two and benchmarks several state-of-the-art methods for search clarification. They also propose a pipeline for online and offline evaluation.  Sekulic et al. [82] propose another incremental dataset ClariQ-FKw based on ClariQ. They extracted facet keywords in clarifying questions as the key information. Recently, Yuan et al. [85] release the first multi-modal clarifying question dataset Melon which extends existing study scenarios by incorporating multi-modal information. The construction of the datasets for search clarification in conversational retrieval is crucial and makes the development of the systems possible.
@@ 3.2 Web Search Clarification
The aforementioned studies for search clarification in conversational retrieval are limited by only selecting the clarifying questions and evaluating the response in a fixed set. Different from them, Zamani et al. [86] propose clarification in web search, In web search, the form of the user’s query is more flexible, which can be a conversational query or a keyword query. When the user’s query is ambiguous or faceted, the system can ask the user a clarification question, and provide multiple candidate facets for the user to choose from, which form a clarification pane as shown in Fig. 5 (b). In the clarification pane, the clarification question is usually generated to improve the user’s search experience, and the candidate facets are a set of generated or selected terms representing a group of high-coherency user sub-intents that the user may be interested in. When the user clicks one of the given facets, the query will be reconstructed by concatenating with the selected facet to form a new query and retrieve a new list of web pages [89].
@@ 3.2.3 Generating Candidate Facets
Unlike search clarification in other scenarios, web search clarification not only asks a clarifying question but also provides users with several candidate facets to click on. This facilitates user interaction and is more suitable for open-domain scenarios. Before this, query facets is a common query intent mining solution, aiming to mine comprehensive, multi-dimensional query sub-intents. For example, there are some rule-based and machine learning-based methods to mining these terms from the search result pages of the query [91, 92, 93, 242].
Different from these studies, web search clarification focuses on finding a set of facets that users are most interested in, rather than showing all of them to users, which helps to concentrate on users’ interests with providing better experiences. Following this idea, Hashemi et al. [94, 95] first propose generating these facets using a seq2seq framework.  Samarinas et al. [96] summarize all rule-based, machine learning-based, and pre-trained language model-based methods to generate the candidate facets for search clarification and conclude that BART-based models taking the query and search result snippets as the input show the best performance in almost all evaluation metrics. Zhao et al. [90] deem that candidate facets are usually in the form of lists, so they collect parallel structures from the search result pages and use this information to enhance the generation ability of the BART-based model. Litvinov et al. [97] analyze the whether and how the coherency of candidate facets influence the generation performance and proposed a coherency-based method for generation. Zhao and Dou [98] think that direct generation of candidate facets is not suitable for multi-turn usage, so they propose a framework to generate candidate facets within a multi-turn manner.
Recently, with the development of LLMs, some LLM-based methods have been proposed for generating candidate facets for web search clarification. Among them, Ni et al. [99] conduct a comparative study to compare different training objectives and base models for generating clarification facets. Lee and Kim [100] fully rely on LLMs to generate the facets. We believe that facets generation based on LLMs will become a more effective approach because LLMs contain much general and domain-specific knowledge.
@@ 3.3 Search Clarification for Question Answering
In question answering (QA), the goal is to answer the user’s specific question. However, the question can contain ambiguity, especially under a conversational scenario. Thus, it is necessary to ask a question for disambiguation. In this section, we classify existing studies in search clarification for QA as (1) Community Question Answering and (2) Knowledge-based Question Answering. As shown in Fig. 5 (c), for Community QA, the system usually focuses on ambiguous intents in a domain-specific scenario such as programming, while the clarification usually focuses on some entity ambiguities in the knowledge-based QA.
@@ 3.3.1 Datasets
In some community QA studies [105, 106], the data is extracted from some community QA websites such as StackExchange and Amazon. These websites contain lots of ambiguous questions, user answers, and clarifying questions. Later, Kumar and Black [208] propose a dataset ClarQ based on StackExchange. To our best knowledge, the scale of ClarQ is much larger than other datasets based on StackExchange. On the other hand, in knowledge-based question answering, Abg-CoQA [212] is proposed to tackle ambiguity in conversational QA, which focuses on disambiguation for human-to-human conversations.
@@ 3.3.2 Community Question Answering
Community Question Answering (Community QA) aims for a collaborative online platform where users can ask questions and receive answers from other community members. Building up such platforms relies on collective knowledge beyond a single expert or automated system. In community QA, several users might engage in multiple turns of discussion around a certain topic, thus clarification interactions frequently occur [105]. Therefore, researching how to automatically clarify user questions is crucial. To this end, Braslavski et al. [243] make the first step to analyze clarifying questions in community QA. They study the user questions from two domains and analyze the user behavior and the taxonomy of the asked clarification questions. After that, Rao et al. [105, 106] design different methods to generate good clarifying questions for community QA. Trienes and Balog [107] is the first study investigating whether a community QA question is clear or not. They propose a method of judging whether a question needs further clarification. Kumar et al. [108] study how to rank clarifying question candidates by natural language inference. As the clarification mechanism is complicated in community QA, some studies [244, 245] made some analytical studies about the feature of good clarifying questions, the taxonomy of clarification questions, and the types of user answers. Although these studies have proposed effective methods to generate community QA questions, these methods are difficult to transfer to other open-domain scenarios.
@@ 3.3.3 Knowledge-based Question Answering
In knowledge-based question answering systems, the clarification process usually targets reference issues between entities. Xu et al. [109] propose a dataset CLAQUA focus on solving clarification in knowledge-based question answering. They also figure out single-turn and multi-turn ambiguity and design to tackle these two situations. Nakano et al. [110] explore the impact of generating pseudo-ambiguous and clarifying questions based on sentence structure on the clarification question answering system. This study provides useful guidance for optimizing the clarification question answering system, enabling it to better understand user needs and generate more targeted clarification questions. Kim et al. [111] study using retrieval-augmented large language models to handle ambiguous questions. Lee et al. [22] provide an in-depth analysis for understanding how to interact with users in open-domain question answering systems and how to adjust their behavior according to user needs. These studies are important for understanding users’ ambiguous intent in conversational search by solving the entity ambiguity problem.
@@ 3.4 Domain-specific Search Clarification
Aside from the aforementioned general search clarification scenarios, some studies focus on a closed search scenario, such as Conversational Recommender System, Legal Search, etc, which are also important. Fig. 5 (d) illustrates two close-domain settings, which require the methods to be adapted accordingly to the scenarios. We introduce and summarize related studies for different domain-specific search clarification scenarios in this section.
@@ 3.4.2 Legal Search Clarification
In legal search, users often deliver ambiguous intents due to the complexity of legal documents. Therefore, a clarifying question can be effective in helping the user articulate her complicated search intent. Liu et al. [51] compare the search behavior and outcomes of conversational and traditional search in legal case retrieval. After that, they further explore how buffer mechanisms and generated queries can help improve the performance of conversational agents in legal case retrieval [124]. Furthermore, they also try to study the behavior of conversational agents for retrieving better legal cases. This study provided an in-depth analysis for understanding how conversational agents interact with users and how to adjust their behavior according to user needs. Recently, Liu et al. [125] introduce event schema, where conversational agents can have a better understanding of the user’s potential intent and ask targeted clarifying questions to improve the accuracy and efficiency of retrieving legal cases.
@@ 3.4.3 Search Clarification for Other Scenarios
Search clarification can further be applied in any scenario where the user’s query or question can be ambiguous or faceted. In this section, we list some other scenarios of search clarification as follows which can be used to expand clarification to new scenarios and can provide inspiration: interactive classification [86], image guessing [126], product description [127, 128], collaborative building [129], programming [130, 131], user simulation [132], social or moral situation [13], task-oriented dialogues [23], math [133], spatial reasoning [134], and collaborative dialogue [135].
@@ 3.6 Future Discussion
Search Clarification for LLMs. In Section 3.5, we summarize recent studies about search clarification with LLMs. These studies focus on utilizing the ability of LLMs as a tool for improving search clarification effectiveness in conversational search systems. However, in some cases, the LLM itself can also be seen as a conversational search system that has a large amount of user logs. Thus, search clarification for LLMs themselves is also worth studying. Besides, search clarification is crucial for understanding user input and might also be important for clarifying LLMs’ output or behavior, e.g., explain why retrieve this document. As LLMs enable more fluid, integrated conversations, clarification might become an inserted dialogue management strategy, seamlessly integrated within the retrieval and generation flow to maintain mutual understanding and guide the interaction effectively.
Search Clarification for Retrieval-Augmented Generation. Some current hot search scenarios can also introduce clarification for enhancement, such as Retrieval-Augmented Generation (RAG). For example, when the user asks a question including an ambiguous entity, the system can ask a question to let the user disambiguate the query and then do the retrieval process. The system can also identify ambiguous or faceted intents embodied in the question to ask more natural clarifying questions.
Search Clarification in Vertical Domains. We can explore more natural clarification generation in some vertical fields, such as medical consultation, customer service robots, etc. The design could focus more on the semantic differences in different domains.
@@ 4 Conversational Retrieval
First, although recent neural models, e.g., BERT [249], T5 [250], and recently emerged LLMs [251], have shown the advanced ability of sequence modeling, it is still a long-term challenge to fully interpret the multi-turn natural language form conversations. Second, the transitions of search intents will introduce much noise into the historical context, which is irrelevant to the current search behavior [26, 60]. Thus, it is crucial to address how to denoise the conversational context. Third, the training data for conversational retrieval models is limited. Labeling multi-turn interactions is time-consuming for human annotators, and the relevance signals of passages are difficult to gain. Therefore, various levels of data augmentation techniques, such as generating queries, context-passage pairs, and whole conversations have been developed. Lastly, though conversational retrieval models can usually achieve better performance than query reformulation approaches, the lack of interpretability hampers the understanding of dense representations in conversations. To address these challenges, the existing studies of conversational retrieval can be categorized into four groups: (1) conversation modeling, (2) context denoising, (3) data augmentation, and (4) interpretation.
@@ 4.1 Conversation Modeling
Figure 6: An example of the end-to-end learning of the model for conversational dense retrieval with implicit de-noising. Each historical turn would have a certain degree of correlation with the current turn, e.g., irrelevant, partially relevant, or strongly relevant.
@@ 4.2 Context Denoising
Within a conversational session, not all historical query turns are relevant to address the current user’s information needs. Simply concatenating all historical context would inject unnecessary and noisy turns and thus degrade the performance of the system. Thus, accurately identifying the useful historical context of the conversation is a crucial research question. To this end, context denoising techniques aim to mitigate the negative impact of irrelevant or noisy historical context, which can be categorized as implicit or explicit. Implicit context denoising techniques achieve this goal by relying on model architectures (e.g., attention mechanisms) or training strategies (e.g., contrastive learning), which inherently learn to focus on relevant parts of the context. Then, the weight of irrelevant or noisy information would be reduced or ignored during representation learning without explicitly selecting or removing historical turns. On the other hand, explicit context denoising methods explicitly select or remain relevant historical elements.
@@ 4.2.1 Implicit Context Denoising
Some works attempt to perform context denoising implicitly, and a common practice is shown in Fig. 6, aiming to teach the model to implicitly identify to what extent (e.g., attention scores) a historical turn is relevant to the current query. For example, Yu et al. [144] leverages dense retrieval to embed conversational queries and documents into a shared vector space where semantic similarity can be directly measured. It employs a teacher-student framework, which involves using an existing, well-trained dense retriever as a teacher to guide the student model, by inheriting its document encodings and learning to mimic the teacher’s embeddings for queries. This strategy enhances model ability by capturing the necessary context while ignoring unrelated information to improve retrieval accuracy across conversational turns. Mao et al. [26] develop a framework by combining curriculum learning and contrastive learning to progressively guide a conversational query encoder on how to filter out irrelevant information from the conversation history. It involves generating noisy and clean versions of conversational data and using these to train the model by gradually increasing the complexity of the training data. Krasakis et al. [145] aim to address context denoising by leveraging existing ad-hoc search data to understand and filter out noisy context, improving the relevance judgment of retrieved passages without additional fine-tuning on conversational datasets. Specifically, it contextualizes the query embeddings within the conversation while ensuring that only the current query terms are matched with potential answers. More recently, Mo et al. [146] propose QRACDR to first define a better aligned embedding space for the query representation learning and enable to model to learn to align the query representation with this area via rewritten queries and relevance judgments, which obtains a less noisy context representation implicitly. Then, Lupart et al. [147] follow this principle to leverage LLM to obtain further improvement.
@@ 4.3.2 Contrastive Sample Generation
Sufficient supervision signals, either positive or negative, are important for dense retrieval [253] with contrastive learning [254]. For conversation scenarios, additional contrastive samples can be generated from the conversational history. The basic assumption [29] is among the top-ranked historical retrieved results, the ones highly relevant to the current query can be used as pseudo positive samples, while others are less relevant and can serve as hard negatives for training. Besides, Kim et al. [150] identify a “retrieval shortcut” phenomenon, which occurs when models overly rely on partial conversation history, but ignore the current query and results in topic drift. To this end, Mo et al. [29] mines additional hard negatives from the historical turns to mitigate the dependency on retrieval shortcuts by forcing the model to consider both the current question and the entire conversation history.